how do you work out 4 times the square root of 2 plus 6 times the sqaure root of 50 ??
Answer by dabanfield(803)

how do you work out 4 times the square root of 2 plus 6 times the sqaure root of 50 ??
4*sqrt(2) + 6*sqrt(50) =
4*sqrt(2) + 6*sqrt(25*2) =
4*sqrt(2) + 6*sqrt(25)*sqrt(2)=
4*sqrt(2) + 6*5*sqrt(2) =
4*sqrt(2) + 30*sqrt(2) = 34*sqrt(2)